Are numbers elevated to power 0 are all ways equal to 1 and numbers elevated to power -1 equal to 0?
Answer by rfer(16322)

First half of your statement is true but the second half isn't.
x^-1= the reciprocal 1/x
x^2= 1/x^2
5^-1= 1/5